40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 4 16:46:31 2013, yuta, Update, SUS, oplev calibration for PRM PRM_PIT.pngPRM_YAW.png
    Reply  Tue Mar 5 17:09:30 2013, yuta, Update, SUS, oplev calibration for ITMY ITMY_PIT.pngITMY_YAW.png
       Reply  Tue Apr 2 15:58:40 2013, Jenne, Update, SUS, oplev calibration for ITMX, BS opl_itmx.pngopl_bs.png
          Reply  Tue Apr 2 16:13:10 2013, rana, Update, SUS, oplev calibration for ITMX, BS 
             Reply  Tue Apr 2 16:55:34 2013, Jenne, Update, SUS, oplev calibration online for ITMs, BS, PRM comparison_opl.pdf
                Reply  Tue Apr 2 18:19:30 2013, Jenne, Update, SUS, BS, PRM oplev servos improved PRM_servo_matches_BS.pdfprm_bs_optical_levers_comparison.pdf
                Reply  Tue Apr 30 22:40:57 2013, Jenne, Update, SUS, oplev XY-plots reflect new calibration 
                   Reply  Sun Sep 22 22:05:10 2013, rana, Update, SUS, oplev XY-plots reflect new calibration noise.png
                      Reply  Thu Oct 3 01:06:48 2013, rana, jenne, Update, SUS, oplev XY-plots reflect new calibration etmx_ol.png
                         Reply  Thu Oct 3 09:49:49 2013, Steve, Update, SUS, ETMY oplev laser is sick 
                      Reply  Tue Nov 5 19:55:12 2013, rana, Update, SUS, oplev XY-plots reflect new calibration 
Message ID: 8391     Entry time: Tue Apr 2 16:55:34 2013     In reply to: 8390     Reply to this: 8393   8514
Author: Jenne 
Type: Update 
Category: SUS 
Subject: oplev calibration online for ITMs, BS, PRM 

[Jenne, Gabriele]

We have put in a new EPICS input into the SUS library part, just before the OL_PIT and OL_YAW filter banks, so that the IN1 point is calibrated to microradians.  I recompiled all SUS-related models.  The OPTLEV_SERVO screen has been changed, so that you can see the calibration, and enter a value.  The gains have been reduced by a factor reciprocal to the calibration, so the loop gain is the same.

ETMs, SRM and MCs all have "calibration" numbers of 1, so the numbers aren't really calibrated, they're just the same as always.

It looks like the PRM and the BS are moving significantly (factor of ~30) more than the ITMs at a few Hz!  (Y-axis of plot is urad/rtHz)

comparison_opl.pdf

EDIT JCD:  We need to fix up the MEDM QPD indicators, and the OpLev red lights on the watchdog screen, so they match the new numbers.  Also, Rana turned on the output limiters to 2000 for all oplev servos.

ELOG V3.1.3-